OpenStudy (xmoses1):

A solid metal cylinder with a 4-in. radius and a 10-in. altitude is melted and recast into solid right circular cones each with a 1-in. radius and a 2-in. altitude. The number of cones formed is???

OpenStudy (xmoses1):

@radar

OpenStudy (radar):

Step 1. Find volume of the larger cylinder. Step 2. Divide this volume by the volume of the smaller cylinders. What is the formula for volume of cylinder?

OpenStudy (xmoses1):

phi R^2 times Height is the formula

OpenStudy (radar):

\[V=\pi r ^{2}h\] So the volume of the large cylcinder is \[\pi 4^{2}(10)\]The volume of the smaller cylinder is\[\pi 1^{2}2=2\pi \]also in cubic inches.

OpenStudy (xmoses1):

Ok i that makes sense so far

OpenStudy (radar):